# Link Representation Learning for Probabilistic Travel Time Estimation

### v1

@@ IV-C Joint Distribution for Multiple Trips in a Day
With the distribution in Eq. (8), we can learn the representation matrices ($L$ and $H$) and the two parameter vectors ($\bm{w}_{\mu}$ and $\bm{w}_{d}$) by maximum likelihood. The total number of parameters is $|\mathcal{V}|\times(r_{L}+r_{H}+2)$, and the parameters are shared for all trips on the road network. As mentioned, the parameters are designed for a specific time window of the day. Given the dynamic nature of traffic demand, time-varying link representation matrices $\{L^{t},H^{t}\}$ are used to dynamically model the joint probability distribution. Here, we employ a simple method to discretize time for temporal modeling, with one day divided into several time intervals. Within each interval $t$, a separate set of embedding vectors $\{L^{t},H^{t},\bm{w}_{\mu}^{t},\bm{w}_{d}^{t}\}$ is learned. We omit the time-interval index $t$ for the rest of this paper.
The parameter $\bm{\theta}=\{L,H,\bm{w}_{\mu},\bm{w}_{d}\}$ of the model can be learned by maximizing the log-likelihood of all trip travel time observations. The log-likelihood of observing $\bm{\tau}_{i}$ is
$\mathcal{L}(\bm{\theta};\bm{\tau}_{i})\propto-\frac{1}{2}\left[\log\det(\tilde{\Sigma})+\left(\bm{\tau}_{i}-\tilde{\bm{\mu}}\right)^{\top}\tilde{\Sigma}^{-1}\left(\bm{\tau}_{i}-\tilde{\bm{\mu}}\right)\right],$ (9)
where $\tilde{\bm{\mu}}=A_{i}\bm{\mu}$ and $\tilde{\Sigma}=A_{i}\Sigma_{d}A_{i}^{\top}+B_{i}(I_{Q_{i}}\otimes\Sigma_{p})B_{i}^{\top}$. Calculating the log-likelihood requires computing the inverse and the determinant of $\tilde{\Sigma}$ of size $Q_{i}\times Q_{i}$, with time complexity of $\mathcal{O}(Q_{i}^{3})$. In practice, the number of trips $Q_{i}$ is often much larger than $|\mathcal{V}|$, and this becomes computationally infeasible. The issue can be potentially addressed by using a small batch size, which can still support the learning of all parameters. Nevertheless, this still requires the inversion of a matrix that matches the batch size, and employing a small batch size could result in inefficient training. In fact, we can effectively reduce the computational cost by leveraging the Woodbury matrix identity and the companion matrix determinant lemma:
@@ IV-D Data Augmentation for Link Representation Learning
Figure 3: Illurstration for sub-sampling data augmentation.
@@ IV-E Conditional Travel Time Estimation based on Joint Probability Distribution
$\operatorname{Cov}\left(\tau_{i}^{q},\tau_{i}^{q^{\prime}}\right)=A_{i}^{q}\Sigma_{d}A_{i}^{q^{\prime}}=\sum_{l\in T_{q}}\sum_{l^{\prime}\in T_{q^{\prime}}}\Sigma_{d}(l,l^{\prime}),$ (14)
@@
$\begin{bmatrix}\bm{\eta}\\ \bm{\tau}_{o}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\bm{0}\\ A_{o}\bm{\mu},\end{bmatrix},\begin{bmatrix}\Sigma_{d}&\Sigma_{d}A_{o}^{\top}\\ A_{o}\Sigma_{d}&A_{o}\Sigma_{d}A_{o}^{\top}+B_{o}\left(I\otimes\Sigma_{p}\right)B_{o}^{\top}\end{bmatrix}\right),$ (15)
@@
$\Sigma^{*}$ $=\Sigma_{d}-\Sigma_{d}A_{o}^{\top}(A_{o}\Sigma_{d}A_{o}^{\top}+B_{o}\left(I\otimes\Sigma_{p}\right)B_{o}^{\top})^{-1}\Sigma_{d}A_{o}^{\top}.$
This conditional distribution can be also derived using precision matrix (i.e., the inverse of covariance), as $\bm{\eta}\mid\bm{\tau}_{o}=\bm{v}\sim\mathcal{N}\left(\bm{\mu}^{*},({\Lambda}^{*})^{-1}\right)$, where
@@
This result is consistent with the estimator derived in [1]; however, our model explicitly learns the spatial structure of $\Sigma_{d}$ and $\Sigma_{p}$ using neural networks, instead of using oversimplified prior specifications. For the travel time $\tau$ of a queried trip with link incident matrix $A$, we can derive its distribution conditional on observed trip analytically:
@@
The conditional distribution obtained above, $p\left({\tau}\mid\bm{\tau}_{o}\right)$, can produce travel time statistics (mean/variance) of the query trips. For deterministic prediction, we can simply take the mean $A(\bm{\mu}+\bm{\mu}^{*})$ as the point estimate. The training process is summarized in Algorithms 1.
Algorithm 1 ProbTTE Training Process
@@
1:Initialize link embedding network $E\_Net1()$, $E\_Net2()$
@@
3:$L,H=E\_Net1(l),E\_Net2(l)$
4:$A=\text{zeros}(b,|\mathcal{V}|)$
5:$B=\text{zeros}(b,b\times|\mathcal{V}|)$
@@
12:$\bm{\mu}=f_{\mu}({AL})\in\mathbb{R}^{b\times 1}$
13:$D=\log(1+\exp(f_{d}(H)))\in\mathbb{R}^{b\times b}$
14:$\Sigma=AL(AL)^{\top}+B(I_{b}\otimes(HH^{\top}+D))B^{\top}\in\mathbb{R}^{b\times b}$
15:$\mathcal{L}=-\log\mathcal{N}(\hat{\bm{\tau}}\,|\,\bm{\mu},\Sigma)$
@@ V-A Datasets and Baselines
We evaluate ProbTTE on two publicly available GPS trajectory datasets from taxi/ride-hailing services: Chengdu: With 3,186 links and 346,074 samples spanning 6 days, travel times range from 420 to 2880 seconds, with a mean of 786 seconds; Harbin: With 8,497 links and 1,268,139 samples spanning 6 days, travel times range from 420 to 2994 seconds, with a mean of 912 seconds. The code and data are available at https://github.com/ChenXu02/ProbTTE.
Model Chengdu Harbin
RMSE MAE MAPE(%) CPRS RMSE MAE MAPE(%) CPRS
DeepTTE $181.31$ $130.10$ $17.20$ $224.23$ $162.59$ $18.36$
HierETA $155.26$ $111.34$ $14.68$ $187.93$ $136.45$ $15.62$
MulT-TTE $149.77$ $105.63$ $13.89$ $178.39$ $129.81$ $14.86$
DeepGTT $165.17$ $118.68$ $15.65$ $1.46$ $191.23$ $143.97$ $16.41$ $1.56$
GMDNet $151.43$ $107.73$ $13.97$ $1.31$ $176.98$ $128.11$ $14.65$ $1.41$
ProbTTE $134.79$ $95.11$ $12.31$ $1.18$ $156.94$ $112.15$ $12.82$ $1.24$
ProbTTE 131.25 93.73 12.14 1.15 153.27 111.14 12.71 1.22
Improvement $12.37\%$ $11.27\%$ $12.60\%$ $12.21\%$ $13.40\%$ $13.25\%$ $13.24\%$ $13.48\%$
@@
We also include a variant of our model, ProbTTE in the comparison, which removes the conditional estimation component and is used to evaluate the naive predictive performance of our model when conditional information is unavailable.
We randomly select 70% of the dataset as the training set, 15% as the validation set, and 15% as the testing set. In terms of the model setting of ProbTTE, the batch size is $b=64$, and the embedding dimension of the link representation vector $r_{L}=r_{H}=36$. The time discretization coefficient $p=24$. All models achieve optimal performance by training for 100 epochs on the 12th Gen Intel(R) Core(TM) i9-12900K CPU and NVIDIA Tesla V100 GPU. We choose four general evaluation metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), Mean Absolute Percent Error (MAPE), and Continuous Ranked Probability Scores (CRPS):
$\text{RMSE}=\sqrt{\frac{1}{{\left|{{{\rm{\mathcal{D}}}}}\right|}}\sum\nolimits_{i\in{{{\rm{\mathcal{D}}}}}}{{{({{\tilde{\tau}_{i}}-{\hat{\tau}_{i}}})}^{2}}}},$ (17)
$\text{MAE}=\frac{1}{{\left|{{{\rm{\mathcal{D}}}}}\right|}}\sum\nolimits_{i\in{{{\rm{\mathcal{D}}}}}}{\left|{{\tilde{\tau}_{i}}-{\hat{\tau}_{i}}}\right|},$ (18)
$\text{MAPE}=\frac{1}{{\left|{{{\rm{\mathcal{D}}}}}\right|}}\sum\nolimits_{i\in{{{\rm{\mathcal{D}}}}}}{\left|{\frac{{{{\tilde{\tau}_{i}}-{\hat{\tau}_{i}}}}}{{{{\hat{\tau}_{i}}}}}}\right|},$ (19)
[…]
where $F(\tau_{i})$ is the cumulative distribution function of $\tau_{i}$. $\tilde{\tau}_{i}$ is a randomly sampled value from the distribution of $\tau_{i}$ of query trip travel time and $\hat{\tau}_{i}$ is the ground truth.
@@ V-B Model Evaluation and Ablation Study
We summarize the experimental results in Table II. As can be seen, the proposed ProbTTE framework clearly outperforms all baseline models. In the Chengdu dataset, ProbTTE outperforms the deterministic baselines by over 1.75% in MAPE and the probabilistic baselines by over 1.83% in MAPE and over 0.16 in CRPS. Compared to the best baseline, ProbTTE shows a relative average improvement of 12.11%. In the Harbin dataset, ProbTTE outperforms the deterministic baselines by over 2.15% in MAPE and the probabilistic baselines by over 1.94% in MAPE and over 0.19 in CRPS, respectively. Our model’s performance has a relative average improvement of 13.34%. For the variant model ProbTTE, its performance in terms of MAPE is slightly lower than the full model ProbTTE by 0.17% and 0.11% on the Chengdu and Harbin datasets, respectively. However, it still maintains a significant advantage over the baseline models. This indicates that joint probability modeling across multiple trips can effectively learn correlations between trips, thus improving the accuracy of travel time estimation.
To further investigate the reasons for model superiority, we constructed three variants of ProbTTE focusing on Multi-trip Modeling, Data Augmentation, and Time Discretization:
  • ProbTTE-w/o MT: The multi-trip modeling component is removed. Each batch contains only one trip, and the correlations between trips are not modeled.
  • ProbTTE-w/o DA: The data augmentation component is removed. The trips are not subsampled, and only the original trip samples are used to train the model.
  • ProbTTE-w/o TD: The time discretization component is removed. Only one set of parameters is used to learn the link embedding vector for all time periods.
@@
The results of the ablation experiments are shown in Table III. We can see that multi-trip joint modeling is the primary source of advantage for our model, contributing to a 2.11% and 2.59% improvement in reducing MAPE. The data augmentation based on subsampling has reduced MAPE by 0.52% and 0.44%, achieving performance improvement solely through reasonable segmentation of the original data without adding any model parameters. Time discretization resulted in gains of 0.48% and 0.32% for the model.
@@
ProbTTE 131.25/153.27 93.73/111.14 12.14/12.71 1.15/1.23
(a) $\Sigma_{d}$ during 9:00-10:00 AM
(b) $\Sigma_{p}$ during 9:00-10:00 AM
(c) $\Sigma_{d}$ during 9:00-10:00 PM
(d) $\Sigma_{p}$ during 9:00-10:00 PM
(e) Real adjacency matrix
(f) Link correlation heatmap from $\Sigma_{d}$
(g) Link correlation heatmap from $\Sigma_{p}$
(h) Visualization of embedding vectors
(i) Visualization of link locations
Figure 4: Link correlation visualization.

### v2

@@ IV-C Joint Distribution for Multiple Trips
With the distribution in Eq. (8), we can learn the representation matrices ($L$ and $H$) and the four parameter matrices ($\bm{w}_{\Sigma_{d}}$$\bm{w}_{\Sigma_{p}}$$\bm{w}_{\mu}$, and $\bm{w}_{d}$) by maximum likelihood. The total number of parameters is $|\mathcal{V}|\times(r_{L}+r_{H})+2(r_{L}^{2}+r_{H}^{2})$, and the parameters are shared for all trips on the road network. As mentioned, the parameters are designed for a specific time window of the day. Given the dynamic nature of traffic demand, time-varying link representation matrices $\{L^{t},H^{t}\}$ are used to dynamically model the joint probability distribution. Here, we employ a simple method to discretize time for temporal modeling, with one day divided into several time intervals. Within each interval $t$, a separate set of embedding vectors $\{L^{t},H^{t},\bm{w}_{\Sigma_{d}}^{t},\bm{w}_{\Sigma_{p}}^{t},\bm{w}_{\mu}^{t},\bm{w}_{d}^{t}\}$ is learned. These embedding vectors are learned through a single complete training process, without the need for retraining on specific scenarios. We omit the time-interval index $t$ for the rest of this paper.
With historical/completed trips, we adopt an empirical Bayes approach [4] to estimate the parameters $\bm{\theta}=\{L,H,\bm{w}_{\Sigma_{d}},\bm{w}_{\Sigma_{p}},\bm{w}_{\mu},\bm{w}_{d}\}$, by maximizing the marginal likelihood of all trip travel time observations. The posterior log-likelihood of observing $\bm{\tau}_{i}^{*}$ is
[…]
where $\tilde{\bm{\mu}}=A_{i}\bm{\mu}$ and $\tilde{\Sigma}=A_{i}\Sigma_{d}A_{i}^{\top}+B_{i}(I_{Q_{i}}\otimes\Sigma_{p})B_{i}^{\top}$. Using Eq. (9) as the loss function allows us to estimate link representation using the emprical Bayes methods. Calculating the log-likelihood requires computing the inverse and the determinant of $\tilde{\Sigma}$ of size $Q_{i}\times Q_{i}$, with time complexity of $\mathcal{O}(Q_{i}^{3})$. In practice, the number of trips $Q_{i}$ is often much larger than $|\mathcal{V}|$, and this becomes computationally infeasible. The issue can be potentially addressed by using a small batch size, which can still support the learning of all parameters. Nevertheless, this still requires the inversion of a matrix that matches the batch size, and employing a small batch size could result in inefficient training. In fact, we can effectively reduce the computational cost by leveraging the Woodbury matrix identity and the companion matrix determinant lemma:
@@
In the model, we learn a set of embedding vectors separately for day-specific and trip-specific random effects and further construct the mean and covariance through two sets of branching functions ($\varphi_{\mu},\varphi_{\Sigma_{d}}$) and ($\varphi_{\Sigma_{p}},\varphi_{d}$). To enhance the efficiency of model optimization, we introduce an orthogonality regularization constraint to promote differentiation among the branched parameters derived from the same embedding vector ($L$ or $H$). The orthogonal regularization term is formulated as:
$\mathcal{L}_{\text{orth}}(\bm{w})=\left(\frac{\bm{w}_{\mu}\cdot\bm{w}_{\Sigma_{d}}}{\|\bm{w}_{\mu}\|\|\bm{w}_{\Sigma_{d}}\|}\right)^{2}+\left(\frac{\bm{w}_{\Sigma_{p}}\cdot\bm{w}_{d}}{\|\bm{w}_{\Sigma_{p}}\|\|\bm{w}_{d}\|}\right)^{2}.$ (12)
Incorporating the likelihood loss $\mathcal{L}_{pre}$ and orthogonal regularization $\mathcal{L}_{\text{orth}}$, the final loss function can be formulated as:
[…]
$+\alpha\left(\left(\frac{\bm{w}_{\mu}\cdot\bm{w}_{\Sigma_{d}}}{\|\bm{w}_{\mu}\|\|\bm{w}_{\Sigma_{d}}\|}\right)^{2}+\left(\frac{\bm{w}_{\Sigma_{p}}\cdot\bm{w}_{d}}{\|\bm{w}_{\Sigma_{p}}\|\|\bm{w}_{d}\|}\right)^{2}\right),$
where $\alpha$ is a hyperparameter to trade off the learning rate.
@@ IV-D Data Augmentation for Link Representation Learning
Figure 3: Illustration for sub-sampling data augmentation.
@@ IV-E Conditional Travel Time Estimation based on Joint Probability Distribution
$\operatorname{Cov}\left(\tau_{i}^{q},\tau_{i}^{q^{\prime}}\right)=A_{i}^{q}\Sigma_{d}(A_{i}^{q^{\prime}})^{\top}=\sum_{l\in T_{q}}\sum_{l^{\prime}\in T_{q^{\prime}}}\Sigma_{d}(l,l^{\prime}),$ (16)
@@
[…]
@@
$\Sigma^{*}$ $=\Sigma_{d}-\Sigma_{d}A_{o}^{\top}(A_{o}\Sigma_{d}A_{o}^{\top}+B_{o}\left(I\otimes\Sigma_{p}\right)B_{o}^{\top})^{-1}A_{o}\Sigma_{d}.$
This conditional distribution can be also derived using the precision matrix (i.e., the inverse of covariance), as $\bm{\eta}\mid\bm{\tau}_{o}=\bm{v}\sim\mathcal{N}\left(\bm{\mu}^{*},({\Lambda}^{*})^{-1}\right)$, where
@@
This result is consistent with the estimator derived in [30], which uses an oversimplified diagonal prior specification for $\Sigma_{d}$ and $\Sigma_{p}$. These priors will work fine to derive the posterior link travel time distribution in the case where we have a large number of observed trips. However, when the number of observed/used trips is small, the obtained posterior will be inaccurate. On the other hand, our model explicitly learns the spatial structure of $\Sigma_{d}$ and $\Sigma_{p}$ using neural networks, which can provide accurate posteriors when we only use a small number of observed or related trips. For the travel time $\tau$ of a queried trip with link incident matrix $A$, we can derive its distribution conditional on observed trips analytically:
@@
The conditional distribution obtained above, $p\left({\tau}\mid\bm{\tau}_{o}\right)$ is fully specified by the learned link representations together with the observed trip travel time values. We can use it to produce travel time statistics (mean/variance) of any queried trips. For deterministic prediction, we can simply take the mean $A(\bm{\mu}+\bm{\mu}^{*})$ as the point estimate. The training process is summarized in Algorithms 1.
Algorithm 1 ProbETA Training Process
@@
1:Initialize link embedding network E_Net1()E\_Net1(), E_Net2()E\_Net2()
@@
3:L,H=E_Net1(l),E_Net2(l)L,H=E\_Net1(l),E\_Net2(l)
[…]
@@
[…]
@@ V-A Datasets and Baselines
We evaluate ProbETA on two publicly available GPS trajectory datasets from taxi/ride-hailing services: Chengdu: With 3,186 links and 346,074 samples spanning 6 days, travel times range from 420 to 2,880 seconds, with a mean of 786 seconds; Harbin: With 8,497 links and 1,268,139 samples spanning 6 days, travel times range from 420 to 2,994 seconds, with a mean of 912 seconds. The code and data are available at https://github.com/ChenXu02/ProbETA.
@@
We also include a variant of our model, ProbETA in the comparison, which removes the conditional estimation component and is used to evaluate the naive predictive performance of our model when conditional information is unavailable.
For each dataset, we randomly select 70% of the dataset as the training set, 15% as the validation set, and 15% as the testing set. In terms of the model setting of ProbETA, the batch size is $b=64$, and the embedding dimension of the link representation vector $r_{L}=r_{H}=32$. The time discretization coefficient $p=24$. All models achieve optimal performance by training for 100 epochs on the 12th Gen Intel(R) Core(TM) i9-12900K CPU and NVIDIA Tesla V100 GPU. We choose four general evaluation metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), Mean Absolute Percent Error (MAPE), and Continuous Ranked Probability Scores (CRPS).
@@ V-B Model Evaluation and Ablation Study
V-B1 Main results analysis
We summarize the experimental results in Table II. As can be seen, the proposed ProbETA framework clearly outperforms all baseline models. In the Chengdu dataset, ProbETA outperforms the deterministic baselines by over 1.75% in MAPE and the probabilistic baselines by over 1.83% in MAPE and over 0.16 in CRPS. Compared to the best baseline, ProbETA shows a relative average improvement of 12.11%. In the Harbin dataset, ProbETA outperforms the deterministic baselines by over 2.15% in MAPE and the probabilistic baselines by over 1.94% in MAPE and over 0.19 in CRPS, respectively. Our model’s performance has a relative average improvement of 13.34%. For the variant model ProbETA, its performance in terms of MAPE is slightly lower than the full model ProbETA by 0.17% and 0.11% on the Chengdu and Harbin datasets, respectively. However, it still maintains a significant advantage over the baseline models. This indicates that joint probability modeling across multiple trips can effectively capture and learn the underlying correlations between trips, enabling a more comprehensive understanding of travel time dependencies.
Compared to ProbETA, the conditional estimation capability of ProbETA enhances the model’s responsiveness to real-time external factors. By leveraging the learned covariance matrix, the model can effectively capture random and ongoing external influences, propagating the estimation bias of the conditional trip travel time onto the target trip to adjust the target estimation. This dynamic adjustment mechanism enables more accurate travel time predictions, improving the model’s overall precision and adaptability.
Model Chengdu Harbin
RMSE (s) MAE (s) MAPE(%) CRPS RMSE (s) MAE (s) MAPE(%) CRPS
DeepTTE $181.31$ $130.10$ $17.20$ $224.23$ $162.59$ $18.36$
HierETA $155.26$ $111.34$ $14.68$ $187.93$ $136.45$ $15.62$
MulT-TTE $149.77$ $105.63$ $13.89$ $178.39$ $129.81$ $14.86$
DeepGTT $165.17$ $118.68$ $15.65$ $1.46$ $191.23$ $143.97$ $16.41$ $1.56$
GMDNet $151.43$ $107.73$ $13.97$ $1.31$ $176.98$ $128.11$ $14.65$ $1.41$
ProbETA $134.79$ $95.11$ $12.31$ $1.18$ $156.94$ $112.15$ $12.82$ $1.24$
ProbETA 131.25 93.73 12.14 1.15 153.27 111.14 12.71 1.22
Improvement $2.37\% 11$$1.27\%$ $60\% 12.21$$21\%$ $13.40\%$ $13.25\%$ $13.24\%$ $13.48\%$
V-B2 Structural ablation experiments and analysis
To investigate the reasons for model superiority, we constructed three structure variants of ProbETA focusing on Multi-trip Modeling, Data Augmentation, and Time Discretization:
  • ProbETA-w/o MT: The multi-trip modeling component is removed. Each batch contains only one trip, and the correlations between trips are not modeled.
  • ProbETA-w/o DA: The data augmentation component is removed. The trips are not subsampled, and only the original trip samples are used to train the model.
  • ProbETA-w/o TD: The time discretization component is removed. Only one set of parameters is used to learn the link embedding vector for all time periods.
@@ V-B2 Structural ablation experiments and analysis
The results of the ablation experiments are shown in Table III. We can see that multi-trip joint modeling is the primary source of advantage for our model, contributing to a 2.11% and 2.59% improvement in reducing MAPE. This indicates that our assumption regarding the correlation of trip travel times, modeled through a joint probability distribution, has a significant advantage over the traditional independent Gaussian assumption. Under the same conditions, the joint probability distribution captures the characteristics of multiple trips in a more holistic manner, making it more reasonable and accurate than single-trip modeling. The data augmentation based on subsampling has reduced MAPE by 0.52% and 0.44%, achieving performance improvement solely through reasonable segmentation of the original data without adding any model parameters. The sub-sampling data augmentation technique does not require introducing additional external features or modifying the model structure; it simply represents travel time information at a higher resolution, helping the model capture finer-grained correlations to enhance performance. Time discretization resulted in gains of 0.48% and 0.32% for the model. This improvement is intuitive, as a shorter time span corresponds to a more stable external environment, leading to smaller variations in travel time and better prediction performance. The detailed experiments and analysis of the hyperparameters related to these three structures, as well as the model dimension (rank), are presented in Section V-D.
@@
ProbETA 131.25/153.27 93.73/111.14 12.14/12.71 1.15/1.23
V-B3 Multi-trip hierarchical ablation experiments and analysis
Our multi-trip modeling approach has proven its advantages through structural ablation experiments. To further assess the contribution of each level in our hierarchical multi-trip probabilistic model, we conducted a series of hierarchical ablation experiments, systematically evaluating their impact on overall performance. Specifically, we constructed three variant models based on different levels: ProbETA-mean, ProbETA-inter, and ProbETA-intra.
  • ProbETA-mean: Only modeling the mean travel time of the links, replacing inter-trip and intra-trip covariance with the identity matrix, and the loss function degenerates into MSE.
  • ProbETA-inter: Only modeling the mean travel time and inter-trip covariance of the links, while replacing intra-trip covariance with the identity matrix.
  • ProbETA-intra: Only modeling the mean travel time and intra-trip covariance of the links, while replacing inter-trip covariance with the identity matrix.
The results with a mark “” represent the outputs directly from the model, while those without the mark represent the results after conditional estimation (ProbETA-mean and ProbETA-intra do not perform conditional estimation). The results as shown in Table IV
We found that both inter-trip and intra-trip correlation modeling contribute positively to performance, with inter-trip correlations having the most significant impact. Specifically, incorporating inter-trip covariance reduces MAPE by 1.04%, while intra-trip covariance leads to a 0.58% reduction. Additionally, conditional estimation based solely on inter-trip covariance further decreases MAPE by 0.08%. We interpret ProbETA-inter as a joint probabilistic model that adopts a coarser approach to variance modeling compared to ProbETA. ProbETA-intra functions as an independent heteroscedastic probabilistic model. Since ProbETA-intra lacks the ability to capture travel time dependencies as effectively as joint probabilistic models like ProbETA or ProbETA-inter, its performance is relatively limited. Inter-trip and intra-trip covariance exhibit a complementary relationship. Inter-trip covariance primarily focuses on capturing dependencies between different trips, while intra-trip covariance, akin to a learned trip-specific bias, refines the diagonal elements of the inter-trip covariance—representing the variance of each individual trip. Together, they jointly enable accurate modeling of multi-trip covariance.
TABLE IV: Hierarchical ablation experiment on Chengdu dataset.
Model RMSE (s) MAE (s) MAPE(%) CRPS
ProbETA-mean 155.71{155.71} 110.09{110.09} 14.27{14.27} 1.39{1.39}
ProbETA-intra 144.12{144.12} 102.99{102.99} 13.35{13.35} 1.30{1.30}
ProbETA-inter 139.68{139.68} 99.19{99.19} 12.89{12.89} 1.26{1.26}
ProbETA-inter 139.25{139.25} 98.76{98.76} 12.81{12.81} 1.25{1.25}
ProbETA $134.79$ $95.11$ $12.31$ $1.18$
ProbETA 131.25 93.73 12.14 1.15